MICROWAVE AND AB INITIO STUDY OF (CH_{3})_{3}CCN-SO_{3}

\underline{GALEN SEDO}, and KENNETH R. LEOPOLD, Department of Chemistry, University of Minnesota, 207 Pleasant St., SE, Minneapolis, MN 55455.

The microwave spectrum of the partially bound complex (CH_{3})_{3}CCN-SO_{3} has been recorded. The nitrogen-sulfur bond length is 2.34 Å, which is almost exactly half way between that in weakly bound N_{2}-SO_{3} and the more strongly bonded (CH_{3})_{3}N-SO_{3}. A simple Townes and Dailey analysis of the ^{1}^{4}N nuclear quadrupole coupling constant gives a value of about 0.19 e^{−} transferred away from the (CH_{3})_{3}CCN upon complexation. Ab initio calculations at the MP2/aug-cc-pvtz level yield a binding energy relative to free (CH_{3})_{3}CCN and SO_{3} of 11.0 kcal/mol, which is only about a fourth of that of (CH_{3})_{3}N-SO_{3}. As noted previously for the complex HCCCN-SO_{3},^{a} comparison of the N-S bond length with those of a series of SO_{3} adducts indicates that the proton affinity Lewis base is a good predictor of the properties of the complex. Indeed, spectra of this adduct were readily located on the basis of the proton affinity value for (CH_{3})_{3}CCN.

aS. W. Hunt, D. L. Fiacco, and K. R. Leopold, J. Mol. Spectrosc. 212, 213 (2002).